One integer is 8 less than another integer. The product of the lesser integer and -5 is 28 more than the greater integer. What are the integers?
Answer by CubeyThePenguin(3113)

x = y - 8
-5x = y + 28 ----> x = (y+28)/(-5)

y - 8 = (y + 28)/(-5)
-5(y - 8) = y + 28
-5y + 40 = y + 28
-6y = -12
y = 2

solution: (x, y) = (-6, 2)